OpenStudy (anonymous):

Suppose that people's heights (in centimeters) are normally distributed, with a mean of 165 and a standard deviation of 6. We find the heights of 60 people. (a) How many would you expect to be between 157 and 173 cm tall?

OpenStudy (mathmate):

(a) Calculate the z-values of 157 and 173: Z(157)=(157-165)/6=-4/3 Z(173)=(173-165)/6=4/3 Look up normal distribution table/calculator to find the difference p=P(Z<173)-P(Z<157) where p is the probability of finding people in this height range. Multiply p by the number in the sample to get the expected value.